### v1

# DiffGAN: A Test Generation Approach for Differential Testing of Deep Neural Networks

Abstract
Deep Neural Networks (DNNs) are increasingly deployed across a wide range of applications, from image classification to autonomous driving. However, ensuring their reliability remains a challenge and in many situations, alternative models with similar functionality and accuracy levels are available. Traditional accuracy-based evaluations often fail to capture behavioral differences between such models, particularly when testing datasets are limited, making it challenging to select or optimally combine models. Differential testing addresses this limitation by generating test inputs that expose discrepancies in the behavior of DNN models. However, existing differential testing approaches face significant limitations: many rely on access to model internals or are constrained by the availability of seed inputs, limiting their generalizability and effectiveness. In response to these challenges, we propose DiffGAN, a black-box test generation approach for differential testing of DNN models. Our method relies on a Generative Adversarial Network (GAN) and the Non-dominated Sorting Genetic Algorithm II (NSGA-II) to generate diverse and valid triggering inputs that effectively reveal behavioral discrepancies between models. Our method employs two custom fitness functions, one focused on diversity and the other on divergence, to guide the exploration of the GAN input space and identify discrepancies between the models’ outputs. By strategically searching the GAN input space, we show that DiffGAN can effectively generate inputs with specific features that trigger differences in behavior for the models under test.
Unlike traditional white-box methods, DiffGAN does not require access to the internal structure of the models, which makes it applicable to a wider range of situations. We evaluate DiffGAN on a benchmark comprising eight pairs of DNN models trained on two widely used image classification datasets. Our results demonstrate that DiffGAN significantly outperforms a state-of-the-art (SOTA) baseline, generating four times more triggering inputs, with higher diversity and validity, within the same testing budget. Furthermore, we show that the generated input can be used to improve the accuracy of a machine learning-based model selection mechanism, which dynamically selects the best performing model based on input characteristics and can thus be used as a smart model output voting mechanism when using alternative models together.
Index Terms:
Deep Neural Network, Differential Testing, Test Generation, GAN, NSGA-II.
1 Introduction
Deep Neural Networks (DNNs) have become essential and widely used across various application domains such as image classification, natural language processing, and autonomous driving [1, 2, 3]. Despite their success, ensuring the accuracy of these models remains a challenge, for example when alternative models with similar functionality and accuracy levels are available. Traditional accuracy-based evaluations of such models often fail to capture behavioral differences between them, particularly when testing datasets are limited, and thus selecting or optimally combining models is challenging.
These challenges underscores the need for differential testing, an approach that focuses on generating triggering inputs—test inputs specifically designed to expose behavioral discrepancies between DNN models under test. Differential testing becomes particularly crucial in scenarios such as model selection, model compression, regression analysis, and the use of model ensembles. In the case of model selection, organizations often need to choose from several models that have been trained for the same task, possibly with different architectures, training data, or optimization techniques.
Although these models may achieve similar accuracy on a given test dataset, their performance can vary significantly under different operational conditions. Differential testing helps identify these variations by generating inputs that expose where one model might outperform the others, providing deeper insights that go beyond simple accuracy metrics. Similarly, in model compression, large DNN models are often compressed using techniques such as quantization, pruning, or knowledge distillation to make them suitable for deployment in resource-constrained environments, such as mobile devices or embedded systems. Differential testing is essential here to ensure that the compressed models retain the accuracy of the original uncompressed versions. By generating inputs that highlight behavioral differences between the compressed and original models, differential testing allows developers to better understand the trade-offs involved in compression and to ensure that the model’s critical performance characteristics are preserved. For regression analysis, when DNN models are updated or re-trained, differential testing identifies and generates inputs where the new version performs worse than the original or vice versa, ensuring that updates do not compromise the accuracy of the model. Last, when using ensembles of models, a voting mechanism is required and one needs to be able to learn when to trust which model’s output.
However, the differential testing process presents several challenges. First, the input subspace where highly accurate models mispredict is often extremely limited, within an extremely large input space. This makes it difficult to find the specific inputs that will trigger behavioral differences, especially in models designed to perform well across a wide range of scenarios. Efficient and effective techniques are thus required to reveal meaningful such discrepancies. Second, many models are treated as black-boxes, where only the inputs and outputs are visible without access to the internal workings of the model. This lack of access complicates the generation of test inputs since the internal decision-making processes remain hidden. Third, limited testing budgets constitute another constraint, particularly in scenarios where collecting, labeling, and validating test input is time-consuming and resource-intensive. In cases where models are deployed in environments that require expensive simulators or manual validation, testing costs can quickly escalate, further limiting the resources available for comprehensive testing.
Finally, ensuring the validity of the generated inputs in differential testing poses a significant challenge. Invalid inputs can lead to mispredictions and skew models’ performance assessments. Human feedback is often required to verify the contextual relevance and validity of test inputs, adding complexity and making the process more resource intensive. This manual validation step is difficult to scale, further complicating efforts to evaluate true performance discrepancies between models effectively.
Several test generation approaches for the differential testing of DNN models have been proposed in the literature [4, 5, 6, 7, 8, 9, 10], but they often come with key limitations. Many of these methods are white-box [4, 5, 6, 7], requiring access to the internal structure and parameters of the DNN models under test. This restricts their applicability to proprietary models, where such internal access is typically unavailable due to privacy or intellectual property concerns. Some other approaches [6, 7] assume structural similarities between the models under test, limiting their applicability on DNN models of different architecture. Black-box differential testing methods [8, 9, 10], on the other hand, often rely on customized mutation operators applied to seed inputs to generate triggering inputs. However, these methods are highly dependent on the availability and quality of the seed data, restricting the generation process to variations of existing inputs.
In response to these challenges, we propose DiffGAN, a black-box test generation approach tailored for the differential testing of DNN models with similar accuracy levels on a given test dataset. By leveraging Generative Adversarial Networks (GANs) and the Non-dominated Sorting Genetic Algorithm II (NSGA-II), DiffGAN generates diverse and valid triggering inputs that effectively reveal behavioral disagreements between the models under test. Since DiffGAN operates as a black-box method, it does not require access to the internal structure of the models or assume structural similarities between them, making it suitable for testing any pair of DNN models.
To evaluate DiffGAN, we constructed a benchmark consisting of eight pairs of models (16 DNN models) trained on two widely used image classification datasets. We compared its performance to DRfuzz [10], a SOTA differential testing baseline assuming the same testing budget. Our experimental results demonstrate that DiffGAN significantly outperforms the baseline, generating on average four times more triggering inputs. Moreover, these inputs include a greater number of valid images and exhibit higher diversity, enabling a more comprehensive assessment of the models’ behaviors. We further demonstrate an application of the triggering inputs generated by DiffGAN through the development of a machine learning-based selection mechanism. This system uses a machine learning model to dynamically select between two classifiers, predicting which one will perform better based on the input characteristics. Our results show that training the selection model with inputs generated by DiffGAN leads to significantly more accurate selections compared to those trained with inputs from the baseline approach.
The main contributions of our paper are as follows:
  1. 1.
    We propose DiffGAN, a novel, black-box test generation approach that leverages GANs and NSGA-II to find triggering inputs for differential testing of DNN models, enabling an exhaustive analysis of behavioral disagreements between models with similar accuracy levels.
  2. 2.
    We perform a comprehensive evaluation on a benchmark of eight image classification model pairs, showing that DiffGAN significantly outperforms the state-of-the-art baseline in generating more diverse and valid triggering inputs, given the same testing budget.
  3. 3.
    We show that the triggering inputs generated by DiffGAN can be used to train a machine learning-based mechanism to select the most accurate model based on input characteristics, leading to improved results when compared to models trained on data generated by the baseline. Model selection constitutes a very important application of differential testing.
The remainder of this paper is structured as follows. In Section 2, we define the research problem and outline the key assumptions underlying our approach. Section 3 details the methodology and design of our proposed solution. In Section 4, we present the empirical evaluation. In Section 4.3, we report our results. Section 6 discusses potential threats to the validity of our findings. Section 7 reviews related works. Finally, Section 8 concludes our work.
2 Problem Definition
In this paper, we propose a test generation approach for the differential testing of DNN models to reveal behavioral disagreements between DNN models that exhibit similar accuracy levels.
2.1 Motivations
Differential testing is crucial across a variety of practical applications such as model selection, voting systems, model compression, model regression, and detecting model reuse [8, 9, 11, 10, 12]. For example, consider the context of model selection. In practice, organizations often face the challenge of choosing the most suitable model during the training phase, prior to inference time. This decision process involves choosing from multiple DNN models, each one possibly trained by different providers for the same task, using different model architectures or training datasets, and demonstrating similar accuracy on a testing dataset. In such scenario, the ability to generate inputs that clearly indicate differences in model behavior is invaluable to guide model selection.
Differential testing is also crucial for model output selection at inference time, in the context of ensemble learning or voting systems, where multiple DNN models are combined to obtain a better prediction than with individual models. This type of testing involves revealing triggering inputs that can not only help characterize the conditions under which one model may outperform the others but also enhance our understanding of each model’s operational boundaries. By understanding the conditions under which one model fares better, one can determine at inference time which model’s output to trust.
Differential testing is also crucial in the context of model compression, a process where large DNN models are transformed into more compact versions through techniques such as quantization, pruning, and knowledge distillation [13]. This transformation is aimed at deploying these models in resource-constrained environments while preserving as much of their original accuracy as possible. In these scenarios, differential testing is invaluable as it helps identify how the behavior of the compressed model deviates from the original. By generating triggering inputs that expose these differences, developers can better understand the trade-offs involved in the compression process.
Differential testing also plays a pivotal role in identifying model regression [10], which entails detecting discrepancies between iterations of a deep learning model (original versus updated versions). For instance, an update aimed at enhancing the overall accuracy of a DNN model might inadvertently introduce inaccuracies under specific input conditions. To mitigate this, triggering test inputs are essential, to systematically assess changes in accuracy across updates and thus ensure they do not lead to poorer accuracy in critical conditions.
Additionally, differential testing is useful for detecting model reuse [12]. In industries where proprietary models are crucial assets, it is important to verify that deployed models are original. Differential testing can identify if a model presented as new is actually derived from or modified from an existing model, thus protecting intellectual property and ensuring the authenticity of models.
In conclusion, while differential testing has broad applications, our focus in this work is specifically on model selection. Unlike other contexts such as model compression, regression detection, or model reuse, the model selection scenario presents unique challenges. Here, the goal is not to evaluate how a single model evolves or changes, but rather to compare multiple models developed independently for the same task. These models may exhibit similar accuracy levels but differ in training data, structure, and optimization methods. Focusing on model selection allows us to explore how differential testing can expose behavioral differences between DNN models. This focus is critical because, in practice, selecting the most appropriate model, whether after training or at inference time, can directly impact the systems using these models.
2.2 Challenges
Generating triggering inputs to reveal behavioral disagreements between DNN models under test introduces three main challenges. First, uncovering discrepancies in highly-accurate models is inherently difficult due to the very small input subspace leading to mispredictions. Because they are designed to perform well under a wide range of conditions, we require effective and efficient techniques to identify behavioral disagreements. Second, in many real-world scenarios, we do not have access to the internal information of DNN models. Indeed, such models are often proprietary and deployed on servers, accessible only through cloud APIs. This limitation constrains us to using black-box techniques to generate triggering inputs as we only have access to the inputs and outputs of the DNN models. Finally, the testing budget is often limited, particularly when models (1) require extensive manual labor to label and validate triggering inputs, or (2) operate on expensive simulators or physical devices. These simulators usually provide a controlled environment that replicates true operational scenarios, allowing for rigorous testing without the costs and risks associated with testing models deployed on expensive physical devices. Within a realistic testing budget, we must effectively generate triggering inputs that reveal diverse behavioral disagreements between the models under test.
2.3 Assumptions
In this work, we propose a differential testing approach for DNN models considering the following assumptions:
Black-Box Models: We assume that our DNN models are black-box. The internal DNN workings, such as intermediate computations and gradients, are inaccessible. This assumption is crucial, especially in scenarios involving models received from third-party providers or models trained via federated learning, where internal model details are naturally obscured to protect privacy and intellectual property. Moreover, frameworks such as TensorFlow Lite and ONNX Inference, commonly used for model quantization, usually offer only end-to-end inference APIs. These frameworks do not provide access to intermediate model data, emphasizing efficiency in inference [14].
Pairwise Differential Testing: Our test generation approach primarily focuses on the pairwise comparison of DNN models. This assumption means that our analysis of behavioral differences is based on comparisons between two DNN models at a time, allowing for the generation of triggering inputs for a pair of DNN models.
No Access to the Training Data: We assume that we do not have access to the original training datasets of the models under test. This is a common scenario in many real-world situations where models are proprietary or when model testing is provided as a service for example. Consequently, our ability to generate inputs and interpret model behavior is based solely on observations from models’ outputs and the currently available testing data, rather than insights from the training process.
Similar Accuracy Levels: We assume that each pair of models under test yields similar and high accuracy on the currently available testing dataset. Such generation warrants the effective generation of triggering inputs to reveal behavioral disagreement between the DNN models, to better assess their respective capabilities and thus enable their use in the various practical applications mentioned above.
Limited Testing Budget: We assume that our test generation approach is constrained by a limited testing budget, a valid assumption in many real-world situations when (1) DNNs are deployed on physical devices or with costly simulators in the loop, or (2) when the cost of the manual labeling or the validation of triggering inputs is high. In our study, we determine the testing budget based on the feasible execution time for test generation. We thus restrict the number of inputs to generate and analyze. Furthermore, a fixed and identical testing budget allows for a fair evaluation of alternative differential test generation approaches.
2.4 Differences from Adversarial Inputs
Adversarial samples are inputs subtly modified to mislead a specific single model by incorporating nearly imperceptible changes that may lead to incorrect predictions. Conversely, triggering inputs are designed to expose discrepancies between two versions of a model—for example the original and its compressed counterpart—by demonstrating that they produce divergent outcomes when fed with the same input. This distinction is crucial for understanding how models behave after modification or for comparing the behavior of two distinct models designed for the same task. Tian et al. [9] showed that adversarial samples of the original model rarely act as triggering inputs for compressed models. By using FGSM and CW adversarial methods [9] on the MNIST dataset, the authors found that only 18 out of 10,000 adversarial samples triggered the compressed models, supporting recent findings that model compression effectively defends against adversarial attacks. Also, several recent studies have shown that adversarial attacks are generally less effective at causing mispredictions in compressed models, as these models exhibit increased robustness against such input manipulations [15, 16]. We therefore exclude such kinds of inputs in our analysis.
3 Approach: Reformulation as an NSGA-II Search Problem
A central problem in differential testing, particularly when evaluating DNN models that exhibit similar accuracy levels, is to generate triggering inputs that reveal behavioral disagreements between models. The generation and analysis of such inputs are crucial, as they not only reveal the specific conditions under which one model may outperform another but also enhance our understanding of the models’ operational boundaries. This insight is important for extracting the precise operational contexts in which each model is likely to fare best, thereby informing more targeted and effective application of these models in real-world scenarios. Traditional approaches, such as transformation filters and adversarial techniques, often fall short in producing inputs that introduce novel characteristics or maintain realism, thus limiting their effectiveness in the comprehensive DNN assessment[17, 18] and differential testing of DNNs. To address this challenge, we introduce DiffGAN, a differential testing approach for DNN models that leverages the capabilities of GANs and NSGA-II to drive the generation of diverse triggering inputs. The use of a GAN model for this purpose results from its proven efficacy in generating realistic and new test inputs across different DNN testing scenarios [19, 20, 21, 22]. When adequately trained, GANs can produce novel inputs that are distinct from the training data and belong to the same input domain—inputs that can be classified within the existing training labels.
In a typical GAN architecture, two deep neural networks operate in tandem: a generator, with the goal of creating new images from a random latent vector exploring the GAN’s latent space, and a discriminator, whose purpose is to determine whether a given image is produced by the generator or a direct sample from the training dataset. The ultimate goal is for the generator to produce new inputs to ensure realism and compliance with the training distribution. Note that the latent space in the context of GANs refers to a high-dimensional space that encodes a compressed representation of the data on which the GAN is trained [23, 24].
Since the models under test often exhibit similar high accuracy levels, GANs alone might not be effective in generating diverse triggering inputs that reveal behavioural disagreements between highly accurate models. To overcome these limitations, we leverage NSGA-II to effectively explore the large GAN’s latent space, thereby enhancing the generation of diverse triggering inputs. More specifically, our genetic search aims at finding latent vectors that maximize the generation of inputs produced by the GAN that are both diverse and reveal behavioral disagreements between pairs of DNN models under test. We also rely on NSGA-II since it is specifically adapted to our multi-objective search problem. It is widely used in the literature and has shown good performance in solving many search-based problems in DNN testing [25, 26, 27]. Many existing works in the literature have shown the effectiveness of NSGA-II in navigating the complex latent space of GANs and solving multi-objective optimization problems [28, 29, 30]. Unlike algorithms such as MOSA, the Many Objective Sorting Algorithm [31], which approaches each fitness function in isolation, NSGA-II tries to find solutions that offer diverse trade-offs among multiple objectives. More specifically, our search is driven by two objectives: (1) maximizing the difference between the output probability vectors of the inputs to generate triggering inputs, and (2) maximizing the diversity of generated inputs by the GAN model. We illustrate in Figure 1 our test generation approach and describe in the following the different steps of our approach along with the different components of our search algorithm.
Figure 1: Overview of DiffGAN
3.1 GAN Training Dataset
To effectively train our GAN model, we use the testing dataset available for the models under test. Indeed, access to the original training datasets of these models is often restricted in many real-world scenarios, particularly with proprietary models, hence our reliance on testing datasets.
To enrich the diversity of the GAN’s training data, we enhance the testing dataset of the models under test with five realistic image transformations, encompassing operations such as rotation, shifting, shearing, and adjustments to the images’ brightness and blurriness. These transformations are applied to every image in the dataset, following the same procedures and parameters established by Shorten et al. [32], which serve to ensure (1) the validity of the generated inputs and (2) labels preservation by these transformations. Consequently, the augmented dataset feeding into the GAN’s training process is expanded to five times its initial size, due to the application of five distinct image transformations on each original test input. Finally, we merge the original testing dataset with the augmented data to build the final GAN training dataset.
3.2 GAN Model Training
We use a GAN model [23] in our approach which encompasses a generator to produce new images and a discriminator whose role is to distinguish between the generated images and real images from the GAN’s training dataset. There are unique challenges in evaluating generators in GANs, which differ fundamentally from standard DNNs. In the realm of traditional DNNs, models’ predictions across various tasks are compared against a ground truth, which consists of the actual outcomes or labels. However, generative models, such as the generator component of a GAN, follow a different paradigm. Their output is not a label but a new, synthetic image that did not previously exist in the dataset. As a result, traditional metrics based on accuracy against a ground truth are not applicable.
To effectively evaluate the performance of the generator, we used two SOTA metrics designed for evaluating GAN’s performance: the Fréchet Inception Distance (FID) and the Kernel Inception Distance (KID) [33]. The FID metric quantifies the difference between the distribution of GAN-generated images and the distribution of actual images in the training dataset. Note that in the literature [34], FID scores typically range from zero to a few hundred, with lower scores indicating higher quality in image generation, suggesting better GAN performance. The KID measures the similarity between the distributions of real and generated images by computing the maximum mean discrepancy between their feature representations [33]. Unlike FID, KID is calculated using a kernel function, making it more robust to variations in sample size. KID scores range from zero to one, the former reflecting superior quality in the images generated by the model.
These metrics were chosen for their strengths in assessing both the quality and diversity of images generated by GANs. FID is sensitive to differences in image distributions, making it effective for evaluating the overall alignment between generated and real data distributions. On the other hand, KID is robust against varying sample sizes, providing accurate results even when only a few images are available [33]. Together, FID ensures that the generated images closely match the real distribution, while KID maintains robustness across different dataset sizes, making the two metrics complementary in their ability to provide a thorough evaluation of GAN-generated images [33].
We train the GAN model using our training dataset and perform hyperparameter tuning using the Optuna framework  [^1] https://optuna.readthedocs.io/en/stable/, which is known for its efficiency in automating the optimization process for machine learning models [35]. Optuna was selected for its user-friendly interface, capability to efficiently search ML models hyperparameter spaces, and support for a variety of search strategies. We also employ Optuna’s pruner feature for stopping the evaluation of unpromising trials, which allows for expediting the optimization process. Our tuning process targets hyperparameters, learning rates, epochs, and batch sizes, and aims to minimize the FID and KID scores.
For each GAN model trained in our study, we generate a set of 10,000 images by inputting 10,000 randomly chosen latent vectors into the GAN. These generated images are then compared to a randomly selected subset of 10,000 real images from the GAN training dataset. To ensure the reliability of our results and reduce the impact of randomness, this procedure is conducted ten times for each GAN model. We then compute the average FID and KID scores across the ten iterations for each trained model. The trained GAN that obtains the lowest average scores for both FID and KID is selected as the optimal model.
3.3 Generation of Triggering Inputs
After training the GAN, we employ the NSGA-II search algorithm to effectively explore the GAN’s latent space, guiding the search towards the generation of diverse triggering input images. In this approach, a genetic algorithm manipulates and evolves a population of latent vectors through successive crossover and mutation operations. This evolutionary procedure iteratively refines these latent vectors towards the objectives. The selection process is guided by specific fitness functions, which, in our context, prioritize the selection of latent vectors that maximize the generation of diverse triggering input images. In other words, through successive iterations, we expect to increasingly produce test inputs that reveal a high number of diverse behavioral disagreements between the DNN models under test. To properly translate the generation process into a search problem using NSGA-II, we need to define the following elements.
3.3.1 Individuals and Initial Population
In genetic search, the initial population contains individuals consisting of a set of elements called genes. These genes are connected and form an individual also referred to as a solution. In our context, the initial population contains 100 latent vectors, each of size 100, as recommended in the literature [36]. We use random selection to build our initial population of individuals, thereby ensuring a diverse starting point for our genetic search.
3.3.2 Fitness Functions
The main objective of our approach is to generate test inputs that reveal diverse behavioral disagreements between the DNN models under test. To achieve this, our search process is guided by two fitness functions. The first function aims to maximize the difference between the output probability vectors of the DNN models under test when fed with the generated input images. By maximizing the difference between the output probability vectors, we are more likely to trigger more pronounced behavioral disagreements between the DNN models under test. The second function aims to maximize the diversity of the generated inputs since generating redundant inputs is a waste of computational resources, particularly when faced with constraints such as limited testing budgets and high labeling costs for testing data. This strategy ensures that each input contributes to assessing the models’ accuracy and identifying conditions where one model outperforms another. By combining both fitness functions, we aim to generate diverse test input images that maximize the chances of revealing distinct behavioral disagreements between the DNN models under test.
Divergence-based Fitness Function
This fitness function estimates the divergence between the output probability vectors of the models under test, when fed with the same input image generated by the trained GAN. More specifically, the fitness function takes as input a GAN latent vector $Z$ responsible for generating the candidate test input image. We feed the models under test with such image and compute the Euclidean distance between the resulting output probability vectors $Y_{m1}$ and $Y_{m2}$. The goal of maximizing this distance is to increase the probability of selecting latent vectors that lead to the generation of input images revealing significant behavioral disagreements between the models under test.
The Euclidean distance serves as a straightforward measure of the magnitude of discrepancy between model predictions, capturing how far apart the models’ outputs are in the probability space. The choice of Euclidean distance over other metrics, such as KL divergence [37], is motivated by its efficiency in terms of computation cost. However, the use of Euclidean distance or any other distance metrics alone has a limitation: they do not account for the differences in predicted labels across models. This means that while a high Euclidean distance indicates a discrepancy in the output probabilities, it does not necessarily reflect a disagreement in the models’ final predicted labels. This is particularly critical when the primary objective lies in searching for triggering inputs that lead to different label predictions by the models under test.
We add the following example to illustrate the above limitation. Let us consider that for a generated image (image 1), the Euclidean distance calculated between the output probability vectors P1 = [0.05,0.05,0.05,0.85] from the first model under test (Model 1) and Q1 = [0.15,0.15,0.15,0.55] from the second model under test (Model 2) is 0.3. Further, both models predict the same class for image 1. Then, let us consider another image (image 2), where the calculated Euclidean distance between the output probability vectors P2 = [0.1,0.2,0.3,0.4] from Model 1 and Q2 = [0.15,0.25,0.35,0.05] from Model 2 is also 0.3, despite that these models predicting different classes for image 2. We obtain the same Euclidean distances for both images though clearly image 2 is more interesting given our objectives.
Using a standard Euclidean distance in the context of differential testing is therefore not optimal and indicates the need to refine the fitness function, to ensure it not only considers the magnitude of input differences but also changes in the predicted labels, indicating different models predictions. To address this limitation, we modified the fitness function to reflect differences in model predictions. More specifically, given a GAN latent vector, if the models output different labels for a GAN-generated image, the fitness function increments the Euclidean distance by one to always assign a higher score for triggering input images. Otherwise, it returns the Euclidean distance between the two output probability vectors as illustrated in Equation 1:
$\small D(z)=\begin{cases}\sqrt{\sum_{i=1}^{n}(Y(i)_{m1}-Y(i)_{m2})^{2}}+1&\text{if the models output}\\ &\text{different labels}\\ \sqrt{\sum_{i=1}^{n}(Y(i)_{m1}-Y(i)_{m2})^{2}}&\text{otherwise}\end{cases}$ (1)
Where $z$ represents a GAN latent vector, $D(z)$ denotes the divergence of $z$, and $Y_{m1}$ and $Y_{m2}$ are the output probability vectors of the models under test.
To illustrate the practical impact of our modified fitness function, reusing the examples above, Image 1, a triggering input, receives a fitness score of 1.3, whereas Image 2, a non-triggering input, is assigned a score of 0.3. This adjusted fitness function more accurately characterize inputs that are likely to reveal significant behavioral differences between models, prioritizing inputs that not only produce different probability vectors but also result in distinct classifications by the models.
Diversity-based Fitness Function
Given a latent vector, this fitness function computes the contribution of the corresponding GAN-generated image to the diversity of the collected triggering inputs during the search process. To do so, the diversity-based fitness function computes the minimum cosine distance score between the pixel values of this generated image and those of the set of collected triggering inputs during the search.
To reduce the computational complexity of comparing each generated image against all collected triggering inputs at each iteration of the search process, we approximate such distance by applying a clustering algorithm to group similar inputs. Then, we only return the minimum cosine distance between the generated image and the representative triggering inputs for all clusters. Preliminary experiments confirmed that, despite the cost of clustering, significant time could be saved in computing similarity new images, thus leading to generating more triggering inputs. More specifically, at each iteration of the search process, we apply the HDBSCAN [38] clustering algorithm on the set of the collected triggering input images to group similar inputs in one cluster. We select HDBSCAN for its ability to identify clusters of varying densities without requiring the prior specification of the number of clusters. This adaptability makes HDBSCAN particularly suited in our context.
Prior to clustering, we implement UMAP for dimensionality reduction, which assists in forming more distinct clusters [39, 27] at a lower cost. Subsequently, within each cluster, we identify the representative input—specifically, the input that has the maximum probability of association with its corresponding cluster222 https://hdbscan.readthedocs.io/en/latest/soft_clustering_explanation.html. The fitness function then calculates the diversity of the GAN-generated image by measuring the cosine distances between this image and the representatives of all clusters. Finally, the diversity-based fitness function returns the minimum of these cosine distances to quantify the diversity contribution of the image, as illustrated in Equation 2:
$Diversity(z)=\min_{i=1}^{m}{CosineDistance(image_{z},image_{i})}$ (2)
where $z$ is the latent vector, $image_{z}$ is the generated image by GAN given $z$, $image_{i}$ is the representative image of the $i^{th}$ cluster formed by HDBSCAN, and $m$ is the total number of clusters.
Multi-Objective Search
We need to maximize in our search the two aforementioned fitness functions to drive the generation of diverse triggering inputs. This is therefore a multi-objective search problem that can be formalized as follows:
$\begin{split}\max\limits_{z\in\mathbb{P}}\>\>&Fitness(z)=(Divergence(z),Diversity(z))\\ \end{split}$ (3)
where $\mathbb{P}$ is the population of latent vectors, $z$ is a latent vector, function $Fitness:z\xrightarrow{}\mathbf{R}^{2}$ consists of two real-value objective functions $(Divergence(z),Diversity(z))$, and $\mathbf{R}^{2}$ is the objective space of our optimization problem.
3.4 Genetic Operators
We describe below the two genetic operators in our input generation approach.
3.4.1 Crossover
The crossover operator takes as input two parents (i.e., two latent vectors) and generates new offspring by combining the selected parents. Similar to existing works that employ NSGA-II on GAN latent vectors [28, 40], we use the simulated binary crossover (SBX) in our approach as it is suitable for real-coded individuals (i.e., continuous optimization problem) [40, 41, 28]. Unlike traditional binary crossover operators, which are designed for discrete or binary representations, SBX operates directly on the real values of the latent vectors. Let $V_{1}$ and $V_{2}$ be the two selected parents for crossover. We provide next an example of the resulting offspring after applying crossover:
$Offspring_{1}(i)=0.5[(1+\beta)V_{1}(i)+(1-\beta)V_{2}(i)]$ (4)
$Offspring_{2}(i)=0.5[(1-\beta)V_{1}(i)+(1+\beta)V_{2}(i)]$ (5)
where $\beta$ is a random variable controlling the extent of the crossover and $i$ denotes the index of the latent vectors’ components.
3.4.2 Mutation
Our genetic search also employs a mutation operator to further explore the search space. Following existing works that use genetic algorithms in GAN latent space [28, 40], we also adopt polynomial mutation [42] to address the continuous nature of the latent space. Polynomial mutation applies polynomial-distributed noise to the genes forming the selected individuals in our search. By applying such noise to the genes, it allows for the exploration of nearby solutions while still maintaining proximity to the original solution. This is also essential for efficiently navigating the complex and high-dimensional search space of GAN latent vectors. Polynomial mutation helps generate diverse solutions by introducing variability in the population of the search. This diversity is crucial for escaping local optima and finding globally optimal or near-optimal solutions.
3.5 Search Algorithm
1
Input: Initial population of individuals ( $ind$in $P$ ), maximum running time ($T_{max}$), trained GAN, two DNNs under test
Output: triggering inputs $\alpha$
2
[…]
4
[…]
6while Current Time - $T_{start}\leq T_{max}$ do
7
8      $P_{new}\xleftarrow{}\emptyset$
9       $P_{Cross}\xleftarrow{}\emptyset$
10       $P_{Cross}\xleftarrow{}Cross(P,CrossRate)$
11       $P_{new}\xleftarrow{}Mutate(P_{Cross},MutationRate)$
12
13      foreach $ind\in P$ do
14             $\text{image}_{\text{gen}}\xleftarrow{}\text{GAN}(ind)$;
15            $[OP1,OP2]\xleftarrow{}\text{Model1}(\text{image}_{\text{gen}}),\text{Model2}(\text{image}_{\text{gen}})$
16
17            if generated image is triggering then
18                  $\alpha\xleftarrow{}\alpha\cup\{ind\}$
19
20      $P\xleftarrow{}Select(P,P_{new},Fitness)$
21       $gen=gen+1$;
22return $\alpha$
Algorithm 1 High-level NSGA-II algorithm
The main objective of our search algorithm is to maximize the generation of diverse triggering inputs to effectively identify the discrepancies between the DNN models under test. A high-level overview of our NSGA-II search process is outlined in Algorithm 1. Assuming $P$ represents the initial population, consisting of a set of $|P|$ random latent vectors, the algorithm initiates an iterative procedure, executing the following steps in each generation until the maximum execution time ($T$) is reached (lines 3-14). The search process includes the following steps. First, we create a new empty population $P_{new}$ (line 4). Second, based on predefined crossover and mutation rates, we create offspring using crossover ($Cross$) and mutation ($Mutate$) operators, incorporating newly created individuals to $P_{new}$ (lines 5-7). Then, for each individual, we generate the corresponding image with the trained GAN and store the triggering images in an archive (lines 8-12). Next, we calculate the fitness scores of the new population by computing the diversity- and divergence-based fitness scores of each individual in the population. We employ tournament [43] selection ($Select$) to choose individuals for survival (line 13). Lastly, we update the population by incorporating the fittest individuals, proceeding to the next generation (line 14). The search terminates when the maximum execution time ($T$) is reached and returns the archive of generated triggering inputs (line 15).
We set the NSGA-II parameters in our search algorithm as follows. The crossover rate $CrossRate$ is equal to 90% and the mutation rate $MutationRate$ is equal to 10%, given recommended values in the literature [44, 45, 46]. Finally, we used Google Colab and the Pymoo library [47] to implement our genetic search.
3.6 Filtering Invalid Inputs
Despite performing well on standard metrics such as FID and KID, DiffGAN, like any input generation process, might still produce invalid inputs. To address this issue, we developed a two-stage filtration process for invalid triggering inputs, applied after the final generation step. This process incorporates two state-of-the-art filtration techniques to ensure the quality and validity of the generated images [48, 10]. Due to the computational cost associated with the filtration process, it is applied only after the final generation step rather than throughout the generation iterations with NSGA-II, optimizing both the testing budget and output quality.
3.6.1 Discriminator-Based Filtering
The first stage of our filtration process leverages the trained discriminator from our GAN model. We continuously evaluate our GAN’s performance during training to ensure it reaches a high level of performance (in terms of GAN’s evaluation metrics). After training, we use the discriminator to filter out unrealistic or invalid images. We should recall that when properly trained, the discriminator can be highly effective at distinguishing and removing invalid images that do not align with the desired output distribution. Since the discriminator is trained to differentiate between real and generated images, it serves as a straightforward filter, ensuring that only images deemed valid are retained for further analysis.
This approach aligns with findings from the literature [49, 50, 10], which demonstrate the utility of using discriminators not only during the training process but also for post-generation validation. By leveraging the capabilities of our trained discriminator, we ensure a context-specific and reliable validation process that is tailored to the dataset on which the GAN was originally trained.
3.6.2 SSIM-Based Invalidity Detection
As a complementary validation method, we incorporated Structural Similarity Index (SSIM)-based detection. SSIM is a metric that evaluates the similarity between two images by comparing their luminosity, contrast, and structure. In our method, each generated image is compared with its most similar real counterpart from the original training set. If the SSIM score is low, it indicates that the generated image is likely to be invalid due to its significant divergence from real data. A low similarity score indicates significant visual differences from its most similar real image, making it likely to be unrealistic or invalid.
To determine the SSIM threshold for filtering invalid inputs, we generated 1,000 test inputs using the trained GAN. For each generated input, we calculated its SSIM score by comparing it to the most similar image from the original training dataset. After manually labeling each input as either valid or invalid, we analyzed the distribution of SSIM scores for both categories. This procedure was applied to both datasets used in our experiments. We consistently observed that invalid inputs exhibited SSIM scores below 40%. Based on these findings, we selected 40% as the threshold for filtering out invalid inputs.
By integrating discriminator-based filtering with SSIM-based invalidity detection, we developed a two-stage filtration system for enhancing the quality and validity of GAN-generated images. The discriminator provides the first layer of defense against invalid images, while SSIM-based filtration, further refines this process by evaluating the structural similarity between each generated image and its closest counterpart in the original training data. This dual filtration approach significantly reduces the risk of maintaining invalid triggering inputs, leading to more useful DiffGAN outputs.
4 Empirical Evaluation: Questions and Design
4.1 Research Questions
Our empirical evaluation is designed to answer the following research questions, covering the effectiveness of DiffGAN in terms of finding triggering inputs, their diversity and validity, and our ability to guide the training of a model selection mechanism that better learns the conditions under which a model fares best.
RQ1. Do we generate more triggering inputs than the existing baseline approach with the same testing budget?
We aim to study in this research question the effectiveness of DiffGAN in producing more triggering inputs compared to a baseline, namely DRfuzz, within the same testing budget. Our choice of DRfuzz is carefully justified below, along with a description of the technique. The testing budget in our context is defined by a predetermined and fixed execution time. Generating large numbers of triggering inputs that reveal behavioral disagreement between DNN models is essential to accurately learn when each of the models can be expected to fare better. Relying on too few triggering images would not enable DiffGAN to identify the subtle differences between models, especially those with similar, high accuracy. Therefore, our research question aims to evaluate and contrast the efficiency of DiffGAN against DRfuzz in terms of their ability to generate more triggering inputs within time constraints.
RQ2. How do DiffGAN and DRfuzz compare in terms of generating valid triggering inputs within the same testing budget?
Generating large numbers of triggering inputs is not sufficient. To the largest extent possible, they also need to be valid. We, therefore, aim to compare the validity of the triggering inputs produced by DiffGAN and those generated by DRfuzz. The concept of input validity is crucial for DNN testing in general and differential testing in particular since it ensures that the generated inputs are both relevant and meaningful. More specifically, they should be valid and belong to the targeted input domain. A valid input for a DNN model is defined as one that human experts within the domain can recognize and categorize accurately, using labels from the domain [3]. For example, in the context of handwritten digit recognition, an image is considered valid if it can be recognized as a digit by a human expert, despite potential imperfections. We, therefore, aim to compare in this research question the proportions of valid inputs that DiffGAN and DRfuzz generate. This comparison is crucial because generating a large number of triggering inputs is only beneficial if these inputs are to a large extent valid, as defined above.
RQ3. Does DiffGAN find more diverse triggering inputs than DRfuzz?
This research question centers on the diversity of the triggering inputs generated by DiffGAN, compared to those produced by DRfuzz. Diversity in this context refers to the range of conditions under which the inputs cause the DNN models to exhibit behavioral differences. Further, a diverse set of triggering inputs is indicative of the approach’s ability to explore and expose a broad spectrum of potential weaknesses across the models under test. Hence, we aim to assess not only the quantity and validity of triggering inputs but also the breadth of scenarios these inputs represent.
RQ4. Can training of an ML-based model selection mechanism be guided more effectively using the triggering inputs generated by DiffGAN compared to those produced by DRfuzz?
We aim in this research question to determine whether the triggering inputs generated by DiffGAN can more effectively guide the training of a model selection mechanism, allowing it to learn the conditions under which one DNN model outperforms another. In many real-world scenarios, it is common to encounter multiple DNNs that perform similarly in terms of overall accuracy, but exhibit significant variations in their performance under specific input conditions. This presents a challenge for developing a model selection mechanism that can dynamically choose the most suitable model based on the input characteristics without having to run all models in real-time. An effective solution to this problem is to use an ML-based selection mechanism, which can predict which of the available models will perform better for a given input. However, training such an ML-based model selection mechanism requires diverse and informative training data that captures and reveals a wide range of behavioural differences between the models. This is where the generation of triggering inputs becomes critical for the selection mechanism to more effectively learn under which condition one model is preferable over the other. Indeed, the effectiveness of the selection mechanism heavily depends on the quality of the triggering inputs used for training. Therefore, we aim to explore in this RQ whether DiffGAN can better guide the training process compared to DRfuzz. It is important to clarify that we do not intend to propose a novel ensemble method or voting mechanism in this work. Instead, this research question focuses on investigating an important application scenario, that is studying whether DiffGAN helps learn which model fares better under various input conditions and thus train a model selection mechanism.
4.2 Baseline Approach
Since our main focus is to generate triggering inputs that reveal behavioral disagreements between pairs of DNN models, we selected the baseline approaches against which to compare our approach according to the following criteria.
First, the baselines should involve the generation of test inputs for more than one DNN model to reveal differences in DNNs behaviors. This criterion ensures that the baselines under comparison are aligned with our core objective of identifying divergences in DNNs’ behavior.
Second, the baseline approaches should be generally applicable and should not make any assumption on the structural similarity between the DNN models under test (e.g., the approach should not only be applicable to retrained models or quantized ones).
Third, the baselines should be peer reviewed, and open-source implementations of the proposed solutions should be provided. This requirement enables rigorous replicability, thus allowing for a fair comparison of DiffGAN’s effectiveness against established benchmarks.
Finally, we exclude baselines that limit the generation of triggering inputs based on the number of initial test inputs available. Specifically, some approaches cap the maximum number of triggering images to match the size of a seed set—a small subset selected from the initial test dataset. This constraint can introduce bias into comparative evaluations. When methods are assessed within a fixed testing budget, often measured in terms of execution time, such limitations can lead to unfair performance comparisons. If one method is restricted by the number of initial inputs while others are not, it does not provide a level playing field for evaluation. Furthermore, in practical scenarios, the availability of initial test inputs might be limited. This scarcity makes it challenging to gauge a baseline method’s ability to generate additional triggering inputs beyond the initial set. From a practical standpoint, it is also difficult to accurately predict the maximum number of triggering inputs required to uncover all possible behavioral disagreements between the DNNs under test. Without knowing this number in advance, setting a cap based on initial seed inputs could hinder the thoroughness of the testing process.
Based on the first criterion we identified DiffChaser [8], BET [11] Diverget [6], DRfuzz [10], and DFlare [9], as candidates. However, BET and Diverget were excluded due to their lack of open-source implementations. Additionally, DFlare was not selected because its approach to generating triggering inputs is constrained by the number of initial inputs in the testing dataset. The method applies multiple mutation rules to each initial test input to find a triggering input. Once a triggering input is found for a particular initial input, the method stops mutating that input and proceeds to the next one. This process limits the maximum number of possible triggering inputs to the number of initial inputs in the testing dataset, as it generates at most one triggering input per initial input. We also decided not to include DiffChaser in our experiments, as DRfuzz demonstrated superior performance in generating a greater number of trigger inputs [10]. Consequently, we only selected DRfuzz as a baseline to evaluate the effectiveness of our proposed approach.
DRfuzz is an approach designed for generating test inputs that uncover diverse regression faults in updated versions of DNN models. In the context of DNNs, a regression fault occurs when an input, correctly predicted by a previous model version, leads to incorrect predictions in the updated version. DRfuzz initiates this process with seed inputs, upon which it applies various mutation rules to produce new inputs that trigger regression faults. To ensure the validity and relevance of these inputs, DRfuzz employs the discriminator component of a trained GAN to filter invalid inputs. A customized reward function in DRfuzz assigns priority scores to each mutation rule based on their effectiveness in generating diverse triggering inputs from the seed inputs. This selection strategy emphasizes mutation rules that are more likely to uncover behavioral disagreements between model versions. To align DRfuzz with our specific research objectives, we have customized the source code of DRfuzz accordingly. We refined the approach to keep not only the inputs that one model predicts correctly while being mispredicted by the other model, but also any inputs that show behavioral disagreements between the models under test.
4.3 Datasets and Models
We consider different combinations of datasets and models in our empirical evaluation. More specifically, we consider MNIST [51] and Cifar-10 [52], two well-known image recognition datasets. As shown in Figure 2, the MNIST dataset consists of 70,000 black-and-white images depicting handwritten digits, split into a training set of 60,000 images and a test set of 10,000 images. The Cifar-10 dataset comprises 60,000 color images distributed across ten categories (Figure 3), including 50,000 images designated for training and 10,000 images for testing. These datasets were employed in conjunction with 16 different DNN models, which included both custom-developed and widely used architectures such as ResNet, LeNet-5, AlexNet, and inception models [53, 54, 55]. Comprehensive details about the combinations of datasets and models we used, along with their accuracy, are outlined Tables I and II.
Table I: Configurations for MNIST Models
Pair Model Architecture Training Data Optimizer Parameters Accuracy
Pair 1 Model 1 4 Conv, 20 layers MNIST Org (60,000 images) Adam (lr=0.001) #Param: 330,730, Batch: 128, Epoch: 10 99.26%
Model 2 7 Conv, 19 layers MNIST Org (60,000 images) Adam (lr=0.001) #Param:: 327,242, Batch: 128, Epoch: 10 99.50%
Pair 2 Model 3 7 Conv, 33 layers
40000 Org + (10% shifts, 15% zooms/rotations)
on 20,000 MNIST
RMSprop (lr=0.001) #Param: 696,402, Batch: 128, Epoch: 10 99.09%
Model 4 7 Conv, 33 layers
40000 Org+ (5% shifts, 10% zooms/rotations)
on 20,000 MNIST
RMSprop (lr=0.001) #Param: 696,402, Batch: 128, Epoch: 10 99.55%
Pair 3 Model 5 LeNet-5 MNIST Org (60,000 images) SGD (lr=0.01) #Param: 44,426, Batch: 32, Epoch: 10 97.04%
Model 6 LeNet-5 MNIST Org (60,000 images) Adam (lr=0.001) #Param: 44,426, Batch: 32, Epoch: 10 98.93%
Pair 4 Model 7 AlexNet MNIST Org (60,000 images) Adam (lr=0.00001) #Param: 24,733,898, Batch: 128, Epoch: 9 98.85%
Model 8 AlexNet MNIST Org (60,000 images) Adam (lr=0.0001) #Param: 24,733,898, Batch: 256, Epoch: 14 99.33%
Table II: Configurations for Cifar-10 Models
Pair Model Architecture Training Data Optimizer Parameters Accuracy
Pair 1 Model 1 ResNet18 Cifar-10 Org (60,000 images) SGD (lr = 1e-1) #Param: 11183562, Batch=128, Epoch=200 83.38%
Model 2 ResNet20 Cifar-10 Org (60,000 images) SGD (lr = 1e-1) #Param: 274442, Batch=128, Epoch=200 85.71%
Pair 2 Model 3 Inception
40000 Org + (15% shifts, 5%zooms, 10%roations)
on 20,000 Cifar-10
SGD (lr = 1e-1) #Param: 5984858, Batch=128 , Epoch=120 93.02%
Model 4 Inception
40000 Org + (10% shifts, 10%zooms, 15%rotations)
on 20,000 Cifar-10
SGD (lr = 1e-1) #Param: 5984858, Batch=128, Epoch=120 93.27%
Pair 3 Model 5 Inception Cifar-10 Org (60,000 images) SGD (Nesterov=True) #Param: 5984858, Batch=128, Epoch=150 89.10%
Model 6 Inception Cifar-10 Org (60,000 images) Adam (lr = 1e-3) #Param: 5984858, Batch=128, Epoch=150 92.78%
Pair 4 Model 7 ResNet18 Cifar-10 Org (60,000 images) SGD (lr = 1e-1)
#Param: 11183562, Batch=250, Epoch=150
80.85%
Model 8 ResNet18 Cifar-10 Org (60,000 images) SGD (lr = 1e-1)
#Param: 11183562, Batch=128, Epoch=200
83.38%
As shown in the tables, our experimental study spanned four distinct and common sources of variation per dataset across DNN models, including the model structures, training data, optimizers and hyperparameters. Our experiments thus cover 16 different DNN models, each selected to evaluate the effectiveness of our test generation approach. For each model pair, the differences in DNN models are highlighted in bold in Tables I and II. Typical differences across models include:
  1. 1.
    Structural Variations: We considered models with different architectures. For example, in the first configuration in Table I, the two models, although both trained on the MNIST dataset, have a different architecture. More specifically, the first model featuring four convolutional layers among a total of 20 layers, achieved a 99.26% accuracy on the MNIST test dataset. In contrast, the second model, with seven convolutional layers out of a total of 19 layers, achieved a 99.50% accuracy.
  2. 2.
    Training Data Variations: Our analysis also covered models that are architecturally identical but trained on different training datasets. For example, as depicted in the second configuration in Table I, the first DNN model was trained using 40,000 images from the MNIST training dataset and 20,000 MNIST images modified by 10% shifts and 15% zooms/rotations, achieving 99.09% accuracy. Conversely, the second model was trained on 40,000 images from the MNIST training dataset and 20,000 MNIST images adjusted by 5% shifts and 10% zooms/rotations, yielding a slightly higher accuracy of 99.55%.
  3. 3.
    Optimizer Variations: We also considered the scenario where two models are trained with different optimizers. For instance, in a case study involving the Cifar-10 dataset (see Pair 3 in Table II), two inception models were trained using distinct optimizers. Specifically, one model was trained with the SGD optimizer and reached an accuracy of 89.10%, whereas the other model was trained with the Adam optimizer, reaching an accuracy of 92.78%.
  4. 4.
    Hyperparameter Differences: Finally, we considered the scenario where two models are different in terms of hyperparameter values. For instance, we trained two AlexNet models on the MNIST dataset (see Pair 4 in Table I) with varying hyperparameters such as learning rate, epochs, and batch sizes. The first model, set with a learning rate of 0.00001, 9 epochs, and a batch size of 128, reached a 98.85% accuracy. Conversely, the second model, with a learning rate of 0.0001, 14 epochs, and a batch size of 256, demonstrated a 99.33% accuracy.
Figure 2: Examples of original MNIST images (considered valid).
These particular variations were chosen because they represent key aspects of model development and testing that are frequently encountered in practice, for example in frameworks like MLflow [^3] https://mlflow.org to support machine learning development, which is widely used for managing machine learning experiments. MLflow tracks models across these four dimensions, emphasizing their importance in both research and real-world applications.
Further, we selected models ranging from simple architectures, such as LeNet, to more complex ones like AlexNet and ResNet, particularly for the Cifar-10 dataset. This choice allowed us to evaluate our framework across a broad spectrum of models, from smaller, less complex architectures to larger, more sophisticated ones. We could thus assess how well our framework performs under different computational and architectural constraints.
Additionally, we specifically focused on model pairs where the accuracy difference between models was minimal. It ensures that our evaluation captures subtle differences in model performance. Detecting such small variations is crucial for demonstrating that our framework can reliably identify differences even when models achieve nearly identical, high accuracy.
Figure 3: Examples of original Cifar-10 images.
5 Empirical Results
We describe in this section our experimental evaluation and report on the results. We first investigate whether we generate more triggering inputs than the baseline with the same testing budget. We then study the validity and diversity of the generated triggering inputs by both approaches. Finally, we present an important application scenario of the generated triggering inputs, namely the development of a machine learning-based mechanism to predict the best model’s output based on input characteristics, and report the corresponding results.
5.1 RQ1. Do we generate more triggering inputs than the existing baseline approach with the same testing budget?
To answer this research question, we executed DiffGAN and DRfuzz considering the same execution time $T\in[1h,2h,3h]$, across four pairs of DNN models and two distinct datasets, resulting in eight unique combinations (Section 4.3). Because of the randomness in DiffGAN and DRfuzz, we re-executed each of them three times and reported the corresponding number of triggering inputs in Tables III and IV. This decision to limit the execution of the approaches to three iterations was initially due to the considerable execution time associated with the experiments. More specifically, our experiments take 96h (8 model pairs x 6h x 2 approaches) if we execute each approach once. As shown in the tables, DiffGAN consistently outperforms DRfuzz in terms of generating more triggering inputs across all models and datasets. More specifically, our results show that DiffGAN generates on average four times more triggering inputs than DRfuzz when considering the same testing budget. We also calculated the Coefficient of Variation (CV) [56], a measure of relative variability that compares the standard deviation to the mean, independent of the data’s scale. We computed the CV score for each model pair and testing budget across the three runs. A lower CV indicates higher consistency across different runs. For example, in our analysis of the MNIST dataset, DRfuzz showed a CV between 0.33% and 21.23%, while DiffGAN exhibited a CV between 1.33% and 12.20%, suggesting less variability in the number of generated triggering inputs across the different runs. These observations are also consistent for the Cifar-10 dataset. These CV values demonstrate the relative stability and reliability of DiffGAN in generating triggering images, with less variability compared to DRfuzz.
We should note that DRfuzz generates inputs by applying a sequence of mutation rules to seed images, optimizing the order of these transformations to increase the likelihood of finding triggering inputs. However, this approach is by definition limited by the seed images and the specific transformations applied to them, which significantly restricts the potential triggering inputs that can be generated. In contrast, DiffGAN explores a much larger input space by optimizing latent vectors within the GAN’s input space through genetic search. The GAN’s input space inherently offers a broader range of possible image variations than the transformations applied in DRfuzz. By optimizing these latent vectors, DiffGAN can generate novel combinations of inputs that may not be generated through mutation-based methods alone. This broader exploration capability thus allows DiffGAN to produce more triggering inputs, as it can effectively explore the GAN’s latent space to uncover a wider range of inputs that may reveal more behavioral disagreements between the DNN models under test.
Moreover, when comparing the results to the initial set of triggering inputs in the original testing datasets, both DiffGAN and DRfuzz demonstrate significant improvements. The improvement is measured as the ratio of the difference between the number of triggering inputs generated through the differential testing approach and the size of the initial set of triggering inputs, divided by the size of this initial set. Specifically, DiffGAN shows an improvement ranging from 2,660% to 12,727% in the MNIST dataset and from 3,956% to 26,606% in the Cifar-10 dataset. Moreover, DRfuzz achieves an improvement ranging from approximately 954% to 5,465% in configurations with MNIST and from 908% to 3,509% in configurations with Cifar-10. Though both approaches are effective in enhancing the generation of triggering inputs, DiffGAN is significantly more so.
Table III: RQ1 results: Number of generated triggering images for models trained on MNIST dataset using DiffGAN and DRfuzz
Models Configs DRfuzz #TI (Initial) DiffGAN
run1 1229 82 4932
run2 1237 82 5188
1h run3 1228 82 4963
run1 2724 82 8091
run2 2525 82 7539
2h run3 3427 82 6853
run1 3745 82 10077
run2 5056 82 10638
M1, M2 3h run3 4888 82 10839
run1 1368 95 4327
run2 1517 95 4531
1h run3 1702 95 4161
run1 3221 95 5607
run2 3227 95 6980
2h run3 2953 95 5308
run1 4493 95 7143
run2 4737 95 7243
M3, M4 3h run3 5089 95 6880
run1 3743 268 6399
run2 3279 268 8191
1h run3 3044 268 7600
run1 7837 268 12560
run2 7023 268 11911
2h run3 7161 268 10866
run1 10766 268 14927
run2 9601 268 16271
M5, M6 3h run3 10226 268 15430
run1 1478 112 6349
run2 1198 112 5196
1h run3 865 112 5385
run1 2360 112 8195
run2 2088 112 7933
2h run3 2686 112 8046
run1 3246 112 11538
run2 4400 112 13622
M7, M8 3h run3 3141 112 15003
Table IV: RQ1 results: Number of generated triggering images for models trained on Cifar-10 dataset using DiffGAN and DRfuzz
Models Configs DRfuzz #TI (Initial) DiffGAN
run1 36329 1993 80700
run2 36916 1993 80875
1h run3 30284 1993 80952
run1 41345 1993 156002
run2 40009 1993 154942
2h run3 46268 1993 154489
run1 62148 1993 200924
run2 62456 1993 195712
M1, M2 3h run3 66384 1993 199785
run1 7040 618 62596
run2 7203 618 61854
1h run3 6913 618 62609
run1 14725 618 103446
run2 14159 618 104297
2h run3 12262 618 119261
run1 18717 618 156179
run2 20832 618 160324
M3, M4 3h run3 19915 618 178621
run1 9553 959 65544
run2 9851 959 67152
1h run3 9605 959 65544
run1 19248 959 124322
run2 18352 959 126532
2h run3 17120 959 122833
run1 33369 959 224578
run2 34321 959 208223
M5, M6 3h run3 32499 959 208522
run1 26068 1750 74039
run2 26737 1750 73702
1h run3 17765 1750 75850
run1 46005 1750 156716
run2 45730 1750 133703
2h run3 45363 1750 133569
run1 63434 1750 243193
run2 63879 1750 221499
M7, M8 3h run3 62148 1750 218304
Answer to RQ1: DiffGAN significantly and consistently outperforms DRfuzz in generating many more triggering inputs within a given time budget, effectively revealing about four times more behavioral disagreements between DNN models than DRfuzz.
5.2 RQ2. How do DiffGAN and DRfuzz compare in terms of generating valid triggering inputs within the same testing budget?
To systematically compare DiffGAN and DRfuzz in terms of generating valid trigger inputs within a given time budget, we considered the inputs generated in RQ1 by both methods. More specifically, we considered all four model pairs using the MNIST dataset and run both approaches for two hours on these models. As mentioned in RQ1, we ran each method three times to minimize randomness and recorded all images generated by each approach during each run. After completing the runs, for each approach and model pair, we aggregated the images generated from all three runs into a comprehensive dataset. By combining the results from three separate runs for each configuration, we were able to mitigate the effects of randomness in the generation process, thereby increasing the reliability of our comparisons.
Table V: MNIST Improvement
Models Time DRfuzz Imp (%) DiffGAN Imp (%)
M1, M2 1h 1401.62 6030.08
2h 3425.61 9029.67
3h 5465.85 12722
M3, M4 1h 1610.87 4462.81
2h 3103.16 6466.66
3h 4655.79 7095.44
M5, M6 1h 1679.29 3037.69
2h 3494.03 5388.33
3h 2770 4508.1
M7, M8 1h 898.21 4871.43
2h 2291.07 8100.3
3h 3080.65 12990.18
Table VI: Cifar-10 Improvement
Models Time DRfuzz Imp (%) DiffGAN Imp (%)
M1, M2 1h 1631.62 3959.07
2h 2031.51 7559.31
3h 3053.63 10094.88
M3, M4 1h 1054.86 10055.92
2h 2120.68 17201.02
3h 3106.35 23296.37
M5, M6 1h 914.39 6959.54
2h 2409.05 16357.15
3h 3390.65 23611.05
M7, M8 1h 1254.76 4195.18
2h 35699.33 147930
3h 63013.5 219901.5
Removing redundant inputs. To eliminate redundancy from the aggregated dataset, we employed a simple hashing-based technique that focused on the exact byte-level representation of each image. Specifically, we converted each image into its byte form and computed a hash value using Python’s built-in hash() function [^4] Python documentation, hash() function, available at: https://docs.python.org/3/library/functions.html##hash . By comparing these hash values, we identified and removed duplicate images that were identical at the byte level. This method ensured that only unique images—in terms of their pixel data—were retained for further analysis.
Selecting representative samples. Due to the large number of triggering inputs generated by DiffGAN and DRfuzz, we selected a sample of triggering inputs from each aggregated dataset. For each setting, we determined the sample size that can confidently represent the whole corresponding dataset with a 95% confidence level and a margin of error of 5% [^5] As computed from https://www.calculator.net/sample-size-calculator.html. As a result, in our experiments, the sample size varied from 361 to 381 across the four settings that we considered, as shown in Table VII. This approach helped choose a minimal subset that was large enough to provide results that accurately reflected the overall distribution of inputs generated by DiffGAN and DRfuzz [57].
Manual validity checking. From the non-redundant set, we randomly sampled the selected images for validation, according to the sample size determined the previous step . To ensure unbiased evaluation, we shuffled images generated by both DiffGAN and DRfuzz and presented them anonymously to three independent expert evaluators. These experts were tasked with assessing the validity of each image based on predefined criteria. Specifically, they labeled images as either valid or invalid. Invalid images were defined as those that displayed two numbers, no number, or other issues such as (1) Images that were a 50/50 blend of two digits and could not be clearly identified as a specific digit, (2) Incomplete or blended digits that could not be classified into one category by consensus, (3) Images containing random shapes or non-digit drawings, or (4) Images that were completely white or dark, with no discernible content. Examples of valid and invalid inputs generated by both approaches are depicted in Figure 4. To determine the validity of each image, we considered the agreement among the evaluators. If at least two out of three evaluators labeled an image as valid, we considered it valid. To quantify the level of agreement between the evaluators, we employ Cohen’s kappa [58]—a statistical measure that accounts for agreement occurring by chance. In our study, the value of Cohen’s kappa is 70%, indicating a high level of agreement among the evaluators and underscoring the reliability of the manual labeling process.
Bootstrapping. After completing the labeling process, we applied bootstrapping to estimate the probability distribution of the validity ratio of the entire generated inputs. Bootstrapping is a nonparametric, simulation-based technique that allows statistical inference from small samples [59]. By resampling the labeled subset (with replacement) multiple times, this method estimates the probability distribution of the validity ratio of the entire dataset of generated inputs. More specifically, we resampled the labeled subset 1000 times with replacement, calculating the validity ratio for each resample. The resulting distribution of validity ratios provided a more reliable estimate of the dataset’s overall validity, accounting for variability and reducing potential bias from relying on a single sample. This resampling approach allowed us to estimate the confidence interval for the percentage of valid inputs across the entire dataset.
Reporting the total number of valid inputs. We calculated the proportion of valid images for each method by dividing the number of valid images by the total number of sampled images. We then multiplied this validity ratio by the average number of images generated in the runs to estimate the expected validity ratio, the 95% confidence interval, and the number of valid images, as reported in Table VII.
Finally, we compared the expected number of valid inputs generated by DiffGAN and DRfuzz to assess which method was more effective at producing valid triggering inputs within the same time budget of two hours. The results presented in Table VII, clearly suggest that DiffGAN significantly outperforms DRfuzz in generating more valid triggering inputs. For instance, in the first setting, it is estimated that DiffGAN generated 8,292 valid triggering inputs, whereas DRfuzz produced only 1,999 such inputs. This corresponds to a validity ratio of 91.25% for DiffGAN compared to 82.46% for DRfuzz. Importantly, those results were consistent across all the settings we considered. To further analyze the statistical significance of our results we performed a statistical analysis using the Wilcoxon signed-rank test [60], with a significance level of $\alpha=0.05$, to assess whether DiffGAN significantly outperforms DRfuzz in generating valid triggering inputs. The Wilcoxon signed-rank test, a non-parametric method, is used to compare the medians of continuous variables for paired samples and is ideal for our case since it does not require any assumptions about the data distribution.
Instead of considering the number of valid inputs, we relied on the validity ratio due to differences in sample sizes. We evaluated the validity ratios from 4000 samples for DiffGAN and 4000 samples for DRfuzz, selected via bootstrapping across four different scenarios (1000 samples per scenario). The Wilcoxon signed-rank test, with w-statistic = 343465.5 and a p-value of 0.000001, confirmed that DiffGAN significantly outperformed DRfuzz at generating valid triggering inputs across all scenarios.
Table VII: Validity Comparison
Models Method Set Size Av. Validity ratio Validity range (95% conf.) #Est. Valid inputs
M1 and M2 DiffGAN 379 0.9125 [0.8819, 0.9396] 8292
DRfuzz 365 0.8246 [0.7848, 0.8635] 1999
M3 and M4 DiffGAN 377 0.80 [0.76, 0.8329] 6555
DRfuzz 364 0.79 [0.7582, 0.82] 1945
M5 and M6 DiffGAN 381 0.8973 [0.8687, 0.9265] 10286
DRfuzz 377 0.8094 [0.7692, 0.8488] 4882
M7 and M8 DiffGAN 381 0.811 [0.7717, 0.8478] 6535
DRfuzz 378 0.7469 [0.7037, 0.7910] 1430
It is important to note that DiffGAN outperforms DRfuzz in generating valid triggering inputs primarily due to the validation mechanisms it employs throughout the input generation process. While DRfuzz relies solely on the GAN discriminator to assess the validity of generated inputs, DiffGAN incorporates both the generator and the SSIM score, enabling a more effective and comprehensive filtering of invalid inputs. Furthermore, DiffGAN leverages the generative capabilities of GANs, which are specifically designed to produce realistic and diverse inputs, enhancing its ability to generate valid samples. In contrast, DRfuzz applies multiple mutation rules to the same seed image, which can often lead to the generation of more invalid inputs.
Answer to RQ2: DiffGAN consistently and significantly outperforms DRfuzz in generating more valid triggering inputs within the same time budget across all experimental settings.
5.3 RQ3. Does DiffGAN find more diverse triggering inputs than DRfuzz?
To answer this research question, we measured the diversity scores of the generated test input sets for DiffGAN and DRfuzz. We selected well-established diversity metrics that are directly applicable to input sets to measure their diversity scores. Specifically, we used two SOTA metrics: Geometric Diversity [1] and Shannon diversity [61, 62], each leveraging different types of input features.
(a) Valid Images - DRfuzz
(b) Valid Images - DiffGAN
(c) Invalid Images - DRfuzz
(d) Invalid Images - DiffGAN
Figure 4: Examples of Valid and Invalid Images for DRfuzz and DiffGAN
Shannon Diversity, also known as the Shannon-Wiener Index or Shannon Entropy, is a SOTA metric used to quantify diversity in a dataset [63, 62, 64, 65]. It is calculated as follows:
$H=-\sum_{i=1}^{S}p_{i}\log_{b}(p_{i})$ (6)
where:
  • $S$ is the total number of distinct categories (taxa, species, or classes).
  • $p_{i}$ is the proportion of the dataset represented by the $i$-th category.
  • $b$ is the base of the logarithm, which is $e$ (natural logarithm).
The metric computes the sum of the proportions of each category multiplied by the logarithm of those proportions. The negative sign ensures that the index is positive and higher diversity translates into higher values of $H$. To calculate the Shannon diversity for a set of images, each image is first flattened into a one-dimensional array of pixel values. These flattened arrays can then be combined into a single array that represents the distribution of pixel values across the entire image set. The Shannon score, or entropy, is then calculated using this distribution, quantifying the diversity of pixel intensities within the dataset. A higher Shannon score indicates greater diversity within the set of images.
In some contexts, computing the exponential of the Shannon diversity score is used to provide a more intuitive interpretation of the index. Specifically, by calculating $e^{H}$, where $H$ is the Shannon diversity of a set of images, we can derive the effective number of image categories within the set. This value represents the number of equally likely categories that would produce the same level of entropy as observed in the dataset. For example, if $H$ increases from 11 to 12, the exponential value $e^{12}$ is substantially larger (approximately 162,754.79) when compared to $e^{11}$ = 59,874.14, indicating that the dataset is considerably more diverse. This approach allows to translate entropy into a more intuitive measure of diversity, making it easier to understand and compare the diversity of different datasets.
Geometric Diversity is a widely used metric for assessing the diversity of test input sets [1, 66, 67]. To calculate the Geometric Diversity score of an input set, we use features extracted by the VGG16 model. Our previous work [1, 27] has shown that combining the Geometric Diversity metric with VGG16 features is highly effective in accurately capturing and quantifying the diversity of input sets.
The Geometric Diversity metric, denoted as $G(S)$, is computed as follows: Given a dataset $X$ and its feature vectors $V$, the geometric diversity of a subset $S\subseteq X$ is defined by:
$G(S)=\det(V_{s}\times V_{s}^{T})$ (7)
This formula represents the squared volume of the parallelepiped formed by the rows of $V_{s}$, where $V_{s}$ consists of the feature vectors corresponding to the subset $S$. A larger volume indicates greater diversity within the subset $S$ in the feature space.
In this research question, we employed both the Shannon and Geometric diversity metrics to analyze the dataset, aiming to evaluate image diversity from two complementary perspectives. The Shannon diversity metric offers a probabilistic view, emphasizing the distribution of pixel values, while the Geometric diversity metric assesses diversity based on spatial relationships and geometric patterns within the image features extracted by VGG16. By integrating these two metrics, we achieve a more comprehensive understanding of the datasets’ diversity.
To measure the diversity scores of the generated test input sets for DiffGAN and DRfuzz, we used the same subjects and generated triggering inputs as in RQ1. Given that Geometric Diversity scores cannot be normalized based on subset size [1], we ensured a fair comparison by randomly selecting subsets of equal size for both metrics. Specifically, for each approach, subject, and testing budget, we randomly selected 1,000 unique, generated triggering inputs and calculated their Shannon and Geometric Diversity scores. To account for the inherent randomness in DiffGAN and DRfuzz, we repeated this process 30 times and reported the average diversity scores.
The results, as shown in Tables VIII and IX, consistently demonstrate that DiffGAN outperforms DRfuzz in generating more diverse triggering inputs, when considering both metrics, highlighting its effectiveness in exploring a broader range of input variations. Similarly to RQ1, we analysed the variability of our diversity results using the CV score. Our CV analysis shows that both DiffGAN and DRfuzz exhibit low variability across different runs, with DiffGAN demonstrating greater consistency. DRfuzz’s GD diversity has a CV range of 6.76% to 18.44%, while DiffGAN is more stable with a CV range of 2.24% to 14.23%. In terms of Shannon metrics, DRfuzz shows a CV range between 0.09% and 0.70%, whereas DiffGAN is even more consistent, with a CV range of 0.02% to 0.33%. These results underscore the superior stability of DiffGAN’s outputs compared to DRfuzz. Furthermore, the superior performance of DiffGAN in generating more diverse triggering inputs can be attributed to its large latent space, which allows for a greater variety of generated inputs, and its diversity-based fitness function specifically designed to maximize the diversity of generated inputs, a feature that is limited in DRfuzz. Indeed, DRfuzz focuses on triggering behavioral disagreements through limited mutation rules, which, while effective in some contexts, do not inherently prioritize or ensure diversity. As a result, the input sets generated by DRfuzz tend to be less diverse compared to DiffGAN.
Table VIII: Diversity scores of MNIST
DRfuzz DiffGAN
Models Configs GD Exponential Shannon ($e^{H}$) GD Exponential Shannon ($e^{H}$)
M1, M2 1h 422.03 $e^{11.36}=86,361.86$ 530.62 $e^{12.10}=178,632.90$
2h 481.24 $e^{11.34}=84,643.49$ 536.05 $e^{12.16}=191,548.60$
3h 318.04 $e^{11.32}=82,829.03$ 366.21 $e^{12.10}=178,497.79$
M3, M4 1h 438.42 $e^{11.39}=89,882.98$ 727.20 $e^{12.21}=200,195.46$
2h 592.18 $e^{11.39}=89,238.14$ 665.31 $e^{12.27}=213,472.83$
3h 645.69 $e^{11.38}=88,908.85$ 661.64 $e^{12.23}=204,636.72$
M5, M6 1h 410.67 $e^{11.50}=98,610.81$ 712.76 $e^{12.27}=213,402.74$
2h 536.47 $e^{11.43}=92,631.60$ 649.41 $e^{12.23}=204,753.74$
3h 551.06 $e^{11.42}=91,634.01$ 578.24 $e^{12.26}=211,800.43$
M7, M8 1h 116.72 $e^{11.40}=90,159.89$ 481.94 $e^{12.10}=178,751.64$
2h 244.65 $e^{11.40}=90,792.41$ 551.06 $e^{12.12}=183,319.27$
3h 231.87 $e^{11.38}=87,505.95$ 666.11 $e^{12.14}=187,734.83$
Table IX: Diversity scores of Cifar-10
DRfuzz DiffGAN
Models Configs GD Exponential Shannon ($e^{H}$) GD Exponential Shannon ($e^{H}$)
M1, M2 1h 1790.80 $e^{14.71}=2,435,217.87$ 2396.99 $e^{14.84}=2,774,122.83$
2h 2108.90 $e^{14.71}=2,432,731.99$ 2333.51 $e^{14.83}=2,772,404.84$
3h 2318.33 $e^{14.70}=2,430,965.85$ 2457.05 $e^{14.94}=3,073,279.80$
M3, M4 1h 976.64 $e^{14.71}=2,434,515.58$ 2395.81 $e^{14.84}=2,774,819.14$
2h 1117.67 $e^{14.71}=2,432,366.04$ 2339.19 $e^{14.83}=2,772,679.02$
3h 1466.08 $e^{14.70}=2,429,166.95$ 2284.04 $e^{14.83}=2,772,395.01$
M5, M6 1h 711.47 $e^{14.73}=2,453,279.49$ 2387.32 $e^{14.84}=2,774,233.41$
2h 1463.54 $e^{14.71}=2,432,146.96$ 2334.41 $e^{14.83}=2,772,466.97$
3h 2045.84 $e^{14.70}=2,430,669.35$ 2283.65 $e^{14.83}=2,772,148.43$
M7, M8 1h 1893.78 $e^{14.71}=2,432,859.74$ 2378.10 $e^{14.84}=2,774,307.47$
2h 2199.47 $e^{14.69}=2,426,034.24$ 2331.51 $e^{14.83}=2,772,278.57$
3h 2183.13 $e^{14.69}=2,425,208.35$ 2295.15 $e^{14.83}=2,771,688.20$
Figure 5: Accuracy of Random Forest in Selecting the Best Model
Answer to RQ3: DiffGAN consistently outperforms DRfuzz in generating more diverse triggering inputs across all DNN models and datasets.
5.4 RQ4. Can training of an ML-based model selection mechanism be guided more effectively using the triggering inputs generated by DiffGAN compared to those produced by DRfuzz?
We aim in this question to study whether we can better guide the training of an ML-based model selection mechanism that consists of (1) two classifiers (model 1 and model 2) responsible for image classification, and (2) a third model that dynamically selects which of the two classifiers to use based on the input characteristics, which is based on Random Forest (RF). It predicts, based on the characteristics of the incoming input, which of the two classifiers (model 1 or model 2) will produce a more accurate output, without the need to run both models in an online setting. When properly trained with triggering inputs, the RF model should learn when each classifier is more accurate, effectively serving as a decision mechanism for selecting the optimal classifier for a given input. We choose RF because: (1) it scales effectively to handle a large number of features, and (2) its robustness against overfitting [68, 69, 70]. We do not intend to provide a novel voting mechanism here but, instead, provide an example of an important application scenario for the triggering inputs we generate. It also serves as a basis to compare the value of the inputs generated by DiffGAN to enable such model selection and compare it with that of DRfuzz.
To achieve this, we trained the model selection system using two distinct training datasets. The first dataset consists of triggering inputs generated by DiffGAN, while the second contains triggering inputs produced by DRfuzz. To create these datasets, we leveraged the inputs generated by DiffGAN and DRfuzz in RQ1 (with an execution time budget of one hour), considering the four model pairs from the MNIST dataset. For each method and model pair, we randomly sampled 1,200 triggering images, resulting in a total of 9,600 images (1,200 images × 4 model pairs × 2 methods). Given the large volume of unlabeled triggering inputs, we outsourced the classification to a professional labeling company. Each input was labeled with a value from 0 to 9, with an additional label for “invalid” inputs. Invalid images were subsequently filtered out from each of the 1,200-image sets. The number of inputs removed ranged from 21 to 121 for DRfuzz, and from 3 to 34 for DiffGAN. It is important to note that we relied on a single labeling company due to cost and time constraints. This company was chosen for its strong reputation and track record of high-quality labeling. To further ensure the accuracy of the labels, we manually validated the labels of a random sample of 1,000 images. We found that all images were correctly labeled, confirming the reliability of the labeling process.
To compare the accuracy of the RF-based model selection systems, we categorized the images into three groups based on whether one model predicted correctly or if both models failed. We excluded images where both models failed from further analysis due to their low occurrence and lack of relevance to the comparison (i.e., if both models fail, it does not matter which classification model is selected). We trained the RF classifiers using features extracted from a pre-trained VGG16 model, rather than raw pixel data, in an attempt to improve accuracy. The VGG16 model provides feature vectors that capture more abstract and informative patterns than pixel-based features, which is expected to improve the ability of the RF model to distinguish between correct and incorrect predictions. For each method and model pair, we selected 700 valid labeled images and used their VGG16 features to train eight RF-based selection models. To ensure a fair comparison between the RF models trained on different datasets in each scenario, we need a common test dataset. For each scenario, we selected 100 images from the DiffGAN-generated test set and 100 from the DRfuzz-generated test set, creating a combined test set of 200 images. This test set allowed us to evaluate the performance of the RF models using the same test set in an unbiased manner. Additionally, we applied the same feature extraction process as for the training sets.
We report the accuracy levels of each of the four selection models we trained in Figure 5. As clearly visible, the RF-based selection models trained using triggering inputs generated with DiffGAN are far more accurate than those trained with inputs from DRfuzz. For instance, in the first scenario involving two RF-based selection models built using classifiers Model 1 and Model 2 from Table I, the RF model trained with DRfuzz achieved an accuracy of 78%, whereas the one trained with DiffGAN reached a much higher accuracy of 87%. Overall, the accuracy levels for the RF models trained with DiffGAN range between 79% and 87% compared to a range of 70% to 78% for those with DRfuzz inputs. This highlights that DiffGAN produces more effective triggering inputs for training, allowing the RF model to better capture the conditions under which one classifier outperforms the other. As a result, RF models trained using DiffGAN inputs are more accurate in selecting the best-performing classifier, leading to a better model selection mechanism.
This improved accuracy can be attributed to several factors. DiffGAN leverages the combination of GAN and NSGA-II to generate highly diverse triggering inputs. Unlike DRfuzz, which applies mutation rules to existing seed inputs, DiffGAN explores a significantly broader latent input space, enabling the generation of more diverse inputs that expose more pronounced behavioral differences between the models. These behavioral differences are crucial for the RF model to learn the conditions when each classifier performs better. Additionally, DiffGAN optimizes its inputs to both maximize diversity and reveal discrepancies, which also enables the RF model to develop a more accurate understanding of the conditions under which one classifier outperforms the other.
Though the accuracy of RF models can still be improved, our main goal here was to show the usefulness of DiffGAN for one important application, not to optimize model selection. The latter would require the labeling of a much larger dataset and possibly the fine-tuning of the RF models or experimenting with other machine learning techniques, which is out of the scope of this paper.
Answer to RQ4: The ML-based model selection mechanisms trained with DiffGAN triggering inputs are more accurate in selecting the best-performing classifier among alternative models, leading to better model selection mechanisms. This highlights that DiffGAN produces more triggering inputs that are more effective for training, allowing the RF model to better capture the conditions under which one classifier outperforms the other. In turn, this can lead to more accurate voting mechanisms at inference time and thus higher classification accuracy.
6 Threats to Validity
In this section, we discuss the threats to the validity of our study and describe how we mitigated them.
Internal threats to validity concern the causal relationship between the treatment and the outcome. One such threat in our study is the inherent randomness in the experimental setup (RQ1-3), particularly the stochastic nature of DiffGAN and DRfuzz. To mitigate this, we executed each experiment three times, and we observed that the variance across these runs was minimal, leading to consistent conclusions in our RQs. As a result, we concluded that further repetitions, which would come at a high experimental cost, were unnecessary to draw reliable conclusions.
Another potential internal threat is related to the bias in the estimation of the validity of the generated inputs. To address this, we relied on three independent evaluators to assess the validity of the generated inputs manually. By requiring consensus among at least two of the three evaluators for determining validity, we minimized the influence of subjective bias, ensuring a reliable and consistent evaluation process. Additionally, the high level of agreement among evaluators, as measured by Cohen’s Kappa, further confirms the reliability of the validity assessments.
Construct threats to validity concern the relationship between the theory and the observations made. A potential threat to construct validity in our work could arise from the generation of invalid inputs by DiffGAN. To mitigate this, we implemented a filtration process using two SOTA techniques: the SSIM score and the GAN’s discriminator. We experimented with different SSIM score thresholds to find a near-optimal value for filtering out invalid images, manually reviewing all images in a set to verify the adequacy of this threshold. While we acknowledge that this threshold might potentially filter out a few valid images, it effectively removes the majority of invalid inputs. We also recognize that while this process significantly reduces the occurrence of invalid inputs, it does not entirely eliminate them. However, it minimizes their impact, enhancing the overall validity of our results. Another potential threat relates to the construct validity of diversity measurements for the generated input sets. To mitigate this, we used two SOTA diversity metrics: Shannon Diversity and Geometric Diversity, each capturing different aspects of inputs’ variability. The consistent results across both metrics further reinforce our confidence in our diversity assessment of the generated inputs.
External threats to validity concern the generalizability of our study. We mitigate this threat by considering eight different combinations of widely used models and datasets. These combinations feature diverse model pairs varying in architecture, training data, optimizers, and hyperparameters. Additionally, we explored a wide range of testing budgets (execution times) in our experiments and compared our results against a SOTA baseline for the differential testing of DNN models.
7 Related Work
Differential testing approaches for DNN models can be characterized as black-box or white-box, depending on their access requirements to the internals of the DNN models under test. A few black-box differential testing approaches for DNNs have been proposed in the literature.
Xie et al. [8] proposed DiffChaser, a black-box, search-based test generation approach aiming to find behavioral disagreements between DNN models and their quantized versions. They initiate the generation process by selecting seed images. Through iterative mutations and crossovers of these seed images, DiffChaser applies a k-Uncertainty [8] fitness function to guide the generation process, aiming to find discrepancies between a model and its quantized version during testing.
Tian et al. [9] proposed Dflare, a search-based test generation approach that iteratively applies a series of mutation operations to a given seed image until a triggering input is found. They rely on Markov Chains and divergence-based fitness function to guide the selection of mutation operators [9]. The fitness function is designed to prioritize mutated inputs that (1) maximize differences between the output probability vectors of two models or (2) trigger new, previously unobserved probability vectors in the models under test.
You et al. [10] introduced DRfuzz, a differential testing method that detects regression faults in updated versions of DNN models. A regression fault occurs when an input correctly predicted by an earlier model version results in an incorrect prediction post-update. Similar to previous approaches, DRfuzz uses seed inputs and mutation rules to generate new inputs that trigger regression faults. It employs a customized reward function to select mutation rules, favoring those that produce more fault-triggering inputs. DRfuzz also leverages a GAN’s discriminator to filter invalid inputs.
The above test generation approaches heavily depend on the availability and quality of seed data, limiting the process to mutated versions of existing inputs. This dependence significantly narrows the range of test scenarios, as it prevents the introduction of entirely new features (i.e, features that are not present in the initial testing dataset) into the testing process—a capability that DiffGAN notably possesses, enhancing its effectiveness in differential testing by generating diverse and new test input scenarios.
A few white-box differential testing approaches have been proposed as well. DeepExplore [4] is the first white-box differential testing approach for DNN models. It formulates the image generation problem as an optimization problem, which uses gradient-based search techniques to find images that maximize neuron coverage and the number of behavioural disagreements between the DNN models under test.
DeepHunter [5] is a white-box test generation approach for DNNs that uses mutation operators on seed inputs along with coverage metrics to guide the generation of triggering inputs. The approach explores different seed selection strategies, incorporating both diversity-based and recency-based methods, to optimize the seed selection process. DeepHunter supports five different coverage metrics to guide the generation process. However, evaluations on quantized DNNs indicate that no single coverage metric consistently outperforms the others across different quantized DNN models, in terms of generating more triggering inputs.
DiverGet [6] and DeepEvolution[7] are two white-box differential testing approaches for quantized DNN models. They leverage multi-objective search algorithms and domain-specific mutation operators to generate triggering inputs. The generation is guided by two fitness functions. The former fitness function maximizes the divergence in output probability vectors between the models under test, while the latter aims to maximize the coverage of neuron activations within those models. While DeepEvolution is designed to test both original and quantized DNN models for image classification tasks, DiverGet specifically targets DNN models used for hyperspectral image classification [6].
The use of white-box differential testing approaches is hindered by the required access to the internal structure of the DNN models. This can be a significant limitation in practice, particularly when the models under test are proprietary or provided by a third party [1]. Additionally, several white-box approaches [6, 7] rely on comparing neuron activation values or model gradients of DNN models, thus assuming structural similarity among the models being tested. These requirements do not apply to DiffGAN, which, unlike existing white-box differential testing approaches, supports testing across any pair of DNN models without requiring structural similarity or the need for internal access.
8 Conclusion
We presented DiffGAN, a novel black-box test generation framework that leverages GANs and NSGA-II for differential testing of deep neural networks. DiffGAN uses customized fitness functions in NSGA-II to guide the search process, generating diverse and valid triggering inputs that expose behavioral differences between models. To the best of our knowledge, this is the first work addressing black-box differential testing aimed at model selection.
We show that DiffGAN is capable of revealing behavioural differences across models by exploring large input spaces through GAN and NSGA-II’s diversity-driven fitness functions. Our empirical evaluation on eight pairs of image classification models trained on two widely-used datasets demonstrated that DiffGAN significantly outperforms the SOTA differential testing baseline, DRfuzz, by generating on average four times more triggering inputs. These inputs are not only more diverse but also more likely to be valid, allowing for a deeper analysis of the models’ behavior. Furthermore, we showed how these inputs can be applied in a machine learning-based voting mechanism, which dynamically selects the most accurate model based on input characteristics, resulting in improved classification accuracy compared to the baseline approach.
For future work, we aim to explore DiffGAN in dynamic environments to assess the side effects of updating a single ML component on the overall system performance. Rather than comparing pairs of DNN models, we will compare the original system with the updated version. This approach will provide valuable insights into maintaining system reliability and adaptability after updates, particularly in critical areas like autonomous driving, and healthcare.
Acknowledgments
This research was supported by a grant from General Motors, as well as the Canada Research Chair and Discovery Grant programs of the Natural Sciences and Engineering Research Council of Canada (NSERC). Lionel Briand’s contribution was partially funded by the Science Foundation Ireland grant 13/RC/2094-2. We would also like to thank Amirhossein Zolfagharian for his valuable assistance in evaluating the validity of the generated inputs in our experiments.